OpenStudy (anonymous):

Find an equation of the line having the given slope and containing the given point. m=5/6 (8,-4) The equation of the line is y= ?

OpenStudy (luigi0210):

Use \[\LARGE y-y_{1}=m(x-x_{1})\]

OpenStudy (anonymous):

in y=mx+b format or what luigi said

OpenStudy (anonymous):

which one do you need?

OpenStudy (anonymous):

what do yo mean which one do i need? so y=5/6x-4 right?

OpenStudy (anonymous):

you can put it in either form but I prefer y=mx+b

OpenStudy (anonymous):

was i right?